The level of volcanic activity over the last twenty four hours has remained low.
Between 16.00 yesterday and 16.00 today there was 1 hybrid earthquake, 10 rockfall signals and 5 volcano-tectonic earthquakes.
Visual observations of the dome were limited today due to poor visibility. Some new rockfall deposits were observed on the eastern flanks during a helicopter observation flight.
Installation of the site at Harris for continuous GPS occupation was completed today.
Dust levels were slightly raised at all sites except for St. Peter's and St. John's which remained low.
Residents are reminded to remain alert and listen to ZJB radio.